1. The demand is given by P = 20 – 0.1Q, where P is the price and Q is the quantity demanded. The monopolist’s total cost is C = 120 + 2Q + 0.05Q2. Find

a. profit-maximizing price and quantity. (2)

b. deadweight loss in the monopoly profit-maximizing equilibrium. (2)
1
Expert's answer
2020-04-08T09:52:07-0400

"a. TR=P\u00d7Q=20Q-0.1Q^2;"

"MR=MC;"

"MR=(TR)^{'}=20-0.2Q;"

"MC=(TC)'=2+0.1Q;"

"20-0.2Q=2+0.1Q;"

"18=0.3Q;"

"Qm=" "60;"

"Pm=20-0.1\u00d760=" $"14;"

"b." Profit"=TR-TC;"

"TR=P\u00d7Q;"

"TR=60\u00d714=" $"840;"

"TC=120+2\u00d760+0.05\u00d760^2;"

"TC=" $"420;"

Profit"=840-420=" $"420;"